# The Representation and Recall of Interwoven Structured Knowledge in LLMs: A Geometric and Layered Analysis

### v1

@@ 4.2 Attribute geometric interrelationship
5 Relationship in attributes representation: from superposition to separation
The last section explores whether one attribute’s representation can recall related attributes without being mentioned, while this section explores the relationships between attribute representations across layers.
5.1 Attribute representations overlap in intermediate layers but become distinct later
As outlined in Sec.3.1, we trained a linear model for each attribute $A_{j}$ at each layer $l$, yielding a weight vector $\mathbf{w}_{j}^{(l)}$ that represents how attribute $A_{j}$ is stored in the activation space of layer $l$. To analyze attribute relationships across layers, we computed the cosine similarity between weight vectors of different attributes using continuation-style activation sets to minimize language pattern influence.
Fig.9 illustrates the cosine similarity across 80 layers of Meta-Llama-3.1-70B. Notably, in high-dimensional spaces, random vector pairs typically approach orthogonality due to the ‘blessing of dimensionality’. To illustrate this, we randomly sampled vector pairs in an 8129-dimensional space (the activation vector size of Meta-Llama-3.1-70B) and calculated their cosine similarity, with the 99.9% confidence interval (CI) shown in gray. Cosine similarity outside this interval indicates meaningful relationships between attributes. See Appendix C for more details.
Figure 9: Cosine similarity between weight vectors of linear probes for attribute pairs across layers in Meta-Llama-3.1-70B. The shaded area (99.9% CI) represents unrelated boundaries.
5.2 Attribute representations exhibit linear relationships
To directly capture relationships between attributes, we map the representation of $A_{j1}$ to $A_{j2}$ at each layer by training a linear model on the last activation of a fixed prompt template (with PCA reducing the activation dimension to 20), e.g., ‘In the periodic table, the $A_{j}$ of $X_{i}$ is .’ Mapping performance is evaluated by $R^{2}$ scores of 5-fold cross-validation.
Fig.10 shows $R^{2}$ scores across layers for attribute pairs. High scores in early layers reflect token-level differences due to shared prompt template. As layers deepen, $R^{2}$ decreases, indicating a shift to semantic representations. In the intermediate layers, $R^{2}$ rises, revealing attribute overlap, exhibiting linear relationships. In the later layers, $R^{2}$ drops as the model separates attributes for task-specific outputs.
@@ 5.2 Attribute representations exhibit linear relationships
Figure 10: R² scores across layers for mapping attribute pairs in Meta-Llama-3.1-70B using a linear model trained on the last activation of a fixed prompt template.
Both experiments highlight the transition from superposition in the intermediate layers to separation in the later layers. Furthermore, the second experiment demonstrates that even simple linear models are able to capture the relationships between different attributes. The observation also explains why the recall ability of the intermediate layers is stronger, as discussed in Sec.4.
@@ 6 Related work
Intermediate layers matter. Recent studies underscore the importance of intermediate layers in LLMs, emphasizing their role in producing more informative representations for downstream tasks compared to final layers 20, 17, 23, 31. These layers are crucial for encoding abstract knowledge, enabling advanced capabilities like in-context learning and transfer learning, which are vital for understanding and optimizing LLMs 35. Additionally, intermediate layers exhibit distinct patterns of information compression and abstraction, such as reduced entropy, allowing them to efficiently represent complex inputs 7, 34.
Superposition. The Superposition Hypothesis suggests that neural networks can encode far more features than neurons they have by compressing high-dimensional concepts into overlapping, nearly orthogonal representations 2, 29, 15. Instead of assigning features to individual neurons, features are represented as sparse linear combinations across neurons, improving encoding efficiency and reducing interference. Toy models demonstrate that sparsity enhances feature disentanglement, balancing compression and accuracy 8. Early layers encode numerous features with sparse combinations, while intermediate layers focus on higher-level contextual features 11.
Linear representation hypothesis. The linear representation hypothesis suggests that neural networks encode high-level features as linear directions in activation space, enabling easier interpretation and manipulation 26. Probing, introduced by 1, assesses feature encoding in models and builds on findings in word embeddings like GloVe and Word2Vec, which capture semantic relationships through linear structures 24, 28. Empirical support spans various contexts, including spatial and temporal representations 12, sentiment analysis 33, task-specific features 13, and broader relational structures 14.
Non-linear Representations. Although the linear representation hypothesis offers insights into neural network representations, studies have highlighted its limitations and emphasized the significance of non-linear structures. 22 demonstrated that GPT models trained on the game ‘Othello’ required non-linear probes to decode board states. Non-linear structures, such as the ‘pizza’ and ‘clock’ patterns 25, 36, and circular representations observed in tasks like predicting days or months using modular arithmetic prompts 10, reveal the complexity of these representations.
@@ 7 Discussion and conclusions
Building on prior research into how LLMs represent individual entities, our study systematically investigates how LLMs encode and recall interwoven structured knowledge across transformer layers. Intermediate layers emerge as pivotal for encoding factual knowledge and maintaining superimposed representations of related attributes, enabling effective recall of associated features even when not explicitly prompted. In contrast, later layers refine these representations into more distinct and task-specific outputs, prioritizing linguistic coherence. In addition, we uncovered geometric patterns, such as 3D spirals, reflecting relationships like the periodicity in chemical elements. This suggests LLMs encode both linear and non-linear representations aligned with the real geometry of knowledge. The findings provide insights that could inform the development of more interpretable and efficient models, with potential applications in scientific discovery and trustworthy AI systems.
@@
Impact Statement
We believe interpretability in LLMs is essential for AI safety, reducing unintended behaviors and building trust. Understanding how knowledge is stored and recalled across layers can inspire more interpretable, efficient models, advance knowledge editing and scientific discovery.
@@ Acknowledgements
Appendix
Appendix A Attention map detailed results
To investigate how the model prioritizes different parts of the input text, we conducted a preliminary analysis using the 32-layer Meta-Llama-3-8B model. We adopted the attribute $A_{j}$, Period and Group, and iterated over $X_{i}$, consisting of 50 elements, using the prompt template: ‘In the periodic table of elements, the $A_{j}$ of $X_{i}$ is.’ These prompts were input into the language model, and we analyzed the average attention across all attention heads in each transformer layer from the token ‘is’ to all other tokens. The averaged results across different prompts are presented in Fig.A1.
Figure A1: Average attention distribution analysis of the 32-layer Meta-Llama-3-8B model across transformer layers, based on prompts, ‘In the periodic table of elements, the $A_{j}$ of $X_{i}$ is,’ where $A_{j}$ is an attribute (period or group) and $X_{i}$ is an element. The heatmap (left) shows average attention from ‘is’ to all tokens, while line plots (right) depict attention to target tokens (e.g., element and attribute), average attention to other tokens, and attention entropy. Intermediate layers focus on significant token relationships with higher entropy, while later layers (excluding the final layer) show evenly distributed attention and lower entropy, reflecting a shift to broader context integration.
@@ B.1 Layer-wise performance evaluation
Fig.B1 illustrates the prediction error across layers when the activation of the last token across layers is replaced with the predicted activation derived from the geometric space $f(r,g,p)=(r\cos\theta,r\sin\theta,r)$. In the early layers, errors gradually decrease because the model has not yet captured semantic information, and the geometric space is still being constructed. The continuous decline in error reflects the model’s growing ability to capture semantic information and progressively build a coherent geometric representation. By layer 20, the error stabilizes, indicating that these layers effectively encode the periodic and geometric relationships between atomic properties such as atomic number, group, and period.
However, beyond layer 30, the error increases sharply as the model begins outputting non-numeric tokens (replaced with an error value of 50 in the plot, corresponding to the maximum possible error given the 50 elements and atomic numbers). This can be attributed to two factors. First, if the numeric token is not the first output, generating the correct answer requires activations across all token positions. But only the last activation was replaced. Intervening too late disrupts the established flow of activations at other positions, which have already determined the output content. Second, it is also likely due to the model shifting its focus from geometric relationships to higher-level abstractions or context-dependent reasoning in the later layers. Therefore, for the intervention experiments on geometric relationships, we selected layer 20 as it balances effective encoding of geometric relationships and minimizes disruption to the model’s output process.
Figure B1: Variation of Absolute Prediction Errors Across Layers with Intervention. The plot shows the mean absolute error (MAE) for each layer, along with the minimum and maximum error range represented by the shaded region. Missing data points were replaced with a value of 50 before computing the absolute errors.
@@ B.2 Detailed evaluation of geometric spaces
The primary evaluation criterion used in the main text is the absolute error threshold ($\leq 2$), as discussed in detail in Sec.4.2. This metric was chosen because it better captures the accuracy of activation interventions. However, other metrics, such as $R^{2}$, Pearson correlation, and qualitative mapping fidelity, also provide valuable insights. These complementary results are summarized in Table B1.
@@
Table B1: Performance of different low-dimensional spaces for activation intervention. Each space represents a unique pattern, with results assessed using $R^{2}$, Pearson correlation, and percentage of predictions within absolute error $\leq 2$.
@@
In the main paper, we demonstrate two geometric space intervention results; however, other shapes can also be extracted. Fig.B2 shows the extracted linear structure from interventions. While the alignment of points along a straight path indicates the presence of a linear structure, the overlapping points suggest its limitations in distinguishing atomic number. Compared to more expressive shapes like spirals, linear structures may struggle to effectively capture periodic or distinct features.
@@
Figure B2: Linear Structure in the Geometric Space from Intervention Experiments. The figure shows predictions (colored points) and their alignment with the ground truth (gray line). While the linear structure is evident, the overlap of points suggests limitations in capturing distinct element properties.
Appendix C Blessing of dimensionality

### v2

@@ 3 Geometric relationships among attributes
4 Direct attribute recall
In the previous section, we observed that elemental knowledge in LLMs forms a 3D spiral structure. Interestingly, although prompts mentioned only atomic numbers, the embeddings also reflected elemental groups, suggesting that LLMs retrieve both explicit and implicit attributes. To better understand these mechanisms, this section investigates direct attribute knowledge recall and Sec. 5 will explores how LLMs access related but unprompted knowledge.
4.1 From continuity to boundary sharpening
Some elemental attributes, such as group and period, naturally exist in both categorical and numerical forms. This duality enables both classification and regression probing, allowing for direct comparisons that have been underexplored in prior work, which often focused exclusively on a single type.
To examine how LLMs access explicitly mentioned knowledge, we use the last-token residual stream from the continuation style prompt $\mathbf{h}_{j}^{(l)}\in\mathbb{R}^{k}$ as the representation of attribute $A_{j}$, and fit a linear probe to predict its corresponding values via:
$f_{j}^{(l)}(\mathbf{h}_{j})=\mathbf{W}_{j}^{(l)}\mathbf{h}_{j}^{(l)}+\mathbf{b}_{j}^{(l)}$
For categorical attribute forms (e.g., Category, Group, Period), $\mathbf{W}_{j}^{(l)}\in\mathbb{R}^{|C_{j}|\times k}$, $\mathbf{b}_{j}^{(l)}\in\mathbb{R}^{|C_{j}|}$. Predictions are made by:
$\hat{y}^{(l)}=\arg\max_{c\in C_{j}}[f_{j}^{(l)}(\mathbf{h}_{j}^{(l)})]_{c}.$
For continuous attributes, we perform scalar regression by setting $\mathbf{W}_{j}^{(l)}=\mathbf{w}_{j}^{(l)\top}$, $\mathbf{w}_{j}^{(l)}\in\mathbb{R}^{k}$, $\mathbf{b}_{j}^{(l)}\in\mathbb{R}$, yielding:
$\hat{y}^{(l)}=\mathbf{w}_{j}^{(l)\top}\mathbf{h}_{j}^{(l)}+b_{j}^{(l)}$
Probes are trained using 5-fold cross-validation on last-token residual streams. We use a linear Support Vector Machine (SVM) for categorical tasks and Support Vector Regression (SVR) with a linear kernel for continuous tasks. The resulting classification accuracies and regression $R^{2}$ scores are shown in Fig. 5, with best-layer results provided in appendix F.5.
Regression probes reveal that continuous numerical features are effectively represented in intermediate layers, as indicated by high $R^{2}$ values (while not reaching 1, see Appendix F.1). These intermediate layers sometimes even outperform the final layers, suggesting that numerical knowledge is already encoded before the final output stage. This aligns with findings by (Meng et al., 2022), which show that factual knowledge recall is already mediated by intermediate MLP layers.
Figure 5: Linear probing results on last token across layers. (a) Regression ($R^{2}$) for numerical attributes and a random baseline. (b) Classification (accuracy) for categorical attributes. All results use 5-fold cross-validation on last-token residual streams.
In classification probes, intermediate layers perform similarly or even better than final layers for clearly distinct non-numerical categories (e.g., metal vs. non-metal), aligning with prior work (Nanda et al., 2023b). However, they significantly underperform in fine-grained numerical classification—e.g., Period accuracy drops from $\sim\!1.0$ (final) to $\sim\!0.7$, and Group from $\sim\!1.0$ to $\sim\!0.6$.
This suggests that while intermediate layers already encode meaningful numerical structure, additional processing in later layers is required to sharpen boundaries and support accurate fine-grained classification. This aligns with intuition: later layers prepare for discrete token outputs, where clearer classification boundaries must emerge. As shown in Appendix Fig. F.1, the confusion matrix from Layer 40 (70B middle layer) is not perfectly accurate, but most misclassifications fall near the diagonal—further demonstrating that intermediate layers encode coherent numerical structure, albeit with blurred categorical boundaries. These observations may provide useful insights for choosing between intermediate and later-layer embeddings in downstream tasks.
Notably, Llama2 7B shows low accuracy (<0.4) on Group classification compared to Llama3.1 8B (>0.8) (but similar performance in Group regression probing) potentially due to its single number tokenization (splitting numbers like ‘12’ into ‘1’ and ‘2’), which may cause confusion between the representations of output tokens like ‘12’ and ‘1’. In contrast, Llama 3 uses separate tokens for numbers below 1000.
Figure 6: Average $\Delta R^{2}$ across five attributes, with with 95% confidence interval shaded. $\Delta R^{2}=R^{2}_{\text{cont}}-R^{2}_{\text{ques}}$.
4.2 Higher Language Sensitivity in Later Layers
The sharpening of numerical representations into categorical boundaries in later layers suggests that these layers might be shaped by the expected output tokens. This raises a question: does the linguistic structure influence the factual representations across layers?
We compared question-style and continuation-style prompts using linear regression probes. Continuation prompts generally lead to direct generation of fact-related tokens, whereas question-style prompts tend to introduce syntactic fillers (e.g., ‘The’) and are more influenced by surface language patterns.
Fig.6 reports the average delta $R^{2}$ across five attributes, with per-attribute results shown in Fig. F.2 (Appendix). As analyses in earlier sections show stronger semantic signals and higher $R^{2}$ in mid-to-late layers, we focus on depths 0.6–1.0. $\Delta{R^{2}}$ increases in the mid-to-late layers, indicating a growing gap between prompt types. Among the 15 attribute–model combinations (3 models × 5 attributes), 12 show a significant increasing trend (FDR-corrected $p<0.05$), with a median Mann–Kendall $\tau$ of 0.55 (Appendix F.3).
The results indicate that, as depth increases, question prompts become progressively less effective than continuation prompts at encoding factual attributes, hinting that the prompt’s linguistic structure exerts a stronger influence on representations in deeper layers. Interestingly, larger models show a slower increase in $\Delta R^{2}$ across layers than smaller models, suggesting they maintain more stable factual representations across prompt types and thus exhibit a smaller distinction between continuation and question prompts.
The rising $\Delta R^{2}$ suggests that deeper layers increasingly blend factual content with linguistic structure to prepare the final tokens. To further test this, we applied the logit lens (nostalgebraist, 2020) and tuned-lens (Belrose et al., 2023). These analyses estimate the token distribution each layer would produce if decoding were halted at that depth, and show that the correct numerical token becomes highly ranked only in the later layers (Appendix D). Complementary attention statistics (Appendix C) reveal that mid-layers focus tightly on the factual token, whereas later layers spread attention over a wider context—patterns consistent with increased syntactic and contextual integration.
5 Indirect attribute recall
In the previous section, we analyzed direct recall of explicitly mentioned attributes across layers. Our earlier geometric analysis showed that LLMs can also recall related attributes that are not explicitly mentioned. In this section, we explore how related but unmentioned attributes are recalled.
5.1 Middle Layers excel at indirect recall
Figure 7: Average $R^{2}$ scores from regression probing across layers for three prompt types: matching, non-matching, and no-mention. All probes predict a fixed target attribute $A_{j_{1}}$; no-mention uses element token residual streams before any attribute appears. Shaded areas show 95% confidence intervals.
We conducted experiments using linear probing to examine the relationships between distinct attributes. Specifically, we extracted last-token residual streams from continuation prompts that mention attribute $A_{j_{1}}$ (matching) or a different attribute $A_{j_{2}}$ (non-matching). We also extracted the residual stream at the element token position, before any attribute is introduced (no mention). Separate probes were trained for each residual stream dataset, always using labels of attribute $A_{j_{1}}$ as targets. To avoid confounding factors, we selected six attribute pairs without direct linear relationships for non matching probe (see Appendix F.4.1). Average $R^{2}$ curves for all attributes are shown in Fig. 7; detailed case-wise linear probing results appear in Appendix Figs. F.3 and F.4.
Attribute information was detectable across all prompt styles. Intuitively, matching prompts should perform best by providing explicit cues, no-mention comes next as it relies on inference, and non-matching prompts perform worst due to misleading signals. Surprisingly, at intermediate layers (around 0.5 depth), non-matching prompts yielded higher linear $R^{2}$ scores than no-mention prompts, suggesting stronger inter-attribute interactions at these depths. This may reflect entangled representations between related attributes, which we analyze further in Sec. 5.2.
Beyond 60% depth, performance follows the expected trend: matching > no-mention > non-matching. The gap between matching and non-matching prompts increases steadily from 0.6 to 1.0 depth. Across 15 model–attribute tests, 14 exhibited statistically significant divergence (FDR corrected $p<0.05$), with a median Mann–Kendall τ of 0.77 (Appendix Fig.F.5, Table F.3). It suggests that attribute representations become more specialized and context-sensitive in deeper layers. Further analyses in Sec. 5.2 provide a more direct explanation, examining how structural relationships between attributes contribute to this layered specialization.
The fact that the ‘no-mention’ prompts perform best in the early layers may seem counterintuitive; however, this is likely because, unlike the other two scenarios, in the ‘no-mention’ case, the last token is the element itself, which may aid recall. In contrast, matching prompts extract residual streams at the final token (such as ‘is’), requiring holistic semantic understanding. As layer depth increases, semantic clarity improves, enhancing explicitly mentioned attributes and reversing this initial trend.
@@ 5.1 Middle Layers excel at indirect recall
Figure 8: $R_{2}$ scores across layers in Meta-Llama-3.1-70B for linear mappings between attribute pairs using the final residual stream from a fixed prompt.
5.2 Stronger linear correlations in middle layers
To explicitly capture relationships between attribute representation, we train a linear mapping from the representation of attribute $A_{j1}$ to attribute $A_{j2}$ at each model layer. Specifically, we utilize the final residual streams from a fixed prompt template (after applying PCA to reduce the dimensionality to 20). The mapping performance is evaluated using $R^{2}$ scores obtained via 5-fold cross-validation.
Fig. 8 illustrates the variation of $R^{2}$ scores across layers for different attribute pairs. In early layers, $R^{2}$ scores are high; however, this observation alone does not necessarily indicate meaningful attribute-level relationships, as initial representations are predominantly sensitive to token-level similarity (see t-SNE analysis in Fig 1). Due to the use of a fixed input template, the resulting inputs exhibit substantial token-level overlap.
In the intermediate layers, where concept-level understanding is evident (as shown by t-SNE and linear probing), we observe a peak in $R^{2}$ scores. This indicates that even simple linear models can effectively capture relationships between different attributes, reflecting their connection in the learned representation space. This also explains why prompts with non-matching attributes outperform those with no attribute mention at these layers in the last section Sec 5.1. In deeper layers, $R^{2}$ scores decline, suggesting a shift toward specialized representations. Similar conclusions from the linear probing weight analysis further support this, as shown in Appendix E.1.
@@ 6 Related work
Linear representation hypothesis. The linear representation hypothesis suggests that neural networks encode high-level features as linear directions in activation space, enabling easier interpretation and manipulation (Nanda et al., 2023c). Probing, introduced by (Alain, 2016), assesses feature encoding in models and builds on findings in word embeddings like GloVe and Word2Vec, which capture semantic relationships through linear structures (Mikolov et al., 2013; Pennington et al., 2014). Empirical support spans various contexts, including spatial and temporal representations (Gurnee and Tegmark, 2023), sentiment analysis (Tigges et al., 2023), task-specific features (Hendel et al., 2023), and broader relational structures (Hernandez et al., 2023).
Non-linear Representations. Although the linear representation hypothesis offers insights into neural network representations, studies have highlighted its limitations and emphasized the significance of non-linear structures. Non-linear structures, such as the ‘pizza’ and ‘clock’ patterns (Nanda et al., 2023a; Zhong et al., 2024), and circular representations observed in tasks like predicting days or months using modular arithmetic prompts (Engels et al., 2024), reveal the complexity of these representations. Following this, we identify a 3D spiral in element representations aligned with periodic trends, suggesting geometric organization of knowledge in LLMs.
Intermediate layers matter. Recent studies underscore the importance of intermediate layers in LLMs, emphasizing their role in producing more informative representations for downstream tasks compared to final layers (Skean et al., 2025; Kavehzadeh et al., 2024; Ju et al., 2024; Liu et al., 2024). These layers are crucial for encoding abstract knowledge, enabling advanced capabilities like in-context learning and transfer learning, which are vital for understanding and optimizing LLMs (Zhang et al., 2024). Additionally, intermediate layers exhibit distinct patterns of information compression and abstraction, such as reduced entropy, allowing them to efficiently represent complex inputs (Doimo et al., 2024; Yin et al., 2024). Building on these findings, our probe results show that intermediate layers encode knowledge in a continuous form, while sharper categorical boundaries emerge in later layers.
Factual recall. Meng et al. (2022) showed that early MLP layers at the entity token are key to recalling factual associations, while later attention layers propagate this information to the output. Geva et al. (2023) expanded this into a three-stage process—enrichment, transfer, and extraction, revealing subject tokens carry multiple implicit attributes. Nanda et al. (2023b) further validated this through detailed circuit analysis, demonstrating entity token representations linearly encode categorical attributes. Building on this, we investigate how multiple related attributes are jointly represented, and show that intermediate layers encode multiple related attributes in structured and overlapping subspaces, forming geometry-aware representations.
@@ 7 Discussion and conclusions
This study reveals LLMs, despite being trained solely on textual data, develop internal representations that reflect the structured geometry of scientific knowledge. We uncover a 3D spiral structure in LLMs that organizes chemical elements in alignment with the structure of periodic table, suggesting that the model implicitly captures domain-specific regularities without explicit supervision. Through probing experiments, we show that middle layers encode continuous, overlapping attribute subspaces suitable for coarse categorization, while deeper layers sharpen decision boundaries and integrate linguistic structure. Moreover, we find that related attributes are strongly linearly associated in middle layers, enabling indirect recall.
These findings demonstrate that LLMs represent symbolic knowledge not as isolated facts but as geometry-aware, interwoven manifolds. We hope this work inspires further investigation into how LLMs represent and reason about scientific knowledge, such as materials property prediction, and informs the design of downstream embedding-based tasks. We believe interpretability in LLMs is essential for AI safety, reducing unintended behaviors and building trust. Understanding how knowledge is stored and recalled across layers can inspire more interpretable, efficient models, advance knowledge editing and scientific discovery.
@@ Acknowledgements
Appendix A Last token residual stream collection
For each layer $l$, we collect last-token residual streams $\mathbf{h}_{i,j,k}^{(l)}$ from prompts $p_{i,j,k}$ across all elements and templates:
where $f^{(l)}$ denotes the layer-$l$ transformation, $T$ is the token length, and $d$ is the hidden dimension. The initial residual stream $\mathbf{h}_{i,j,k}^{(0)}$ is obtained by embedding the prompt through an embedding layer $E_{0}$, followed by processing through $L$ Transformer layers. Each layer applies multi-head attention and a feedforward network with residual connections and layer normalization:
$\mathbf{h}{\prime}^{(l)}{i,j,k}=\mathbf{h}^{(l-1)}{i,j,k}\;+\;\mathrm{Attention}\!\bigl{(}\mathrm{LayerNorm}(\mathbf{h}^{(l-1)}_{i,j,k})\bigr{)}$
$\mathbf{h}^{(l)}{i,j,k}=\mathbf{h}{\prime}^{(l)}{i,j,k}\;+\;\mathrm{FFN}\!\bigl{(}\mathrm{LayerNorm}(\mathbf{h}{\prime}^{(l)}_{i,j,k})\bigr{)}$
$\text{logits}_{i,j,k}=\mathbf{h}_{i,j,k}^{(L)}\mathbf{W}_{\text{vocab}}$
By analyzing last-token residual streams $\mathbf{h}_{i,j,k}^{(l)}$ across layers, we investigate how attributes are represented in the model’s hidden states.
@@ B.1 Layer-wise performance evaluation
Fig. B.1 illustrates the prediction error across layers when the residual stream of the last token across layers is replaced with the predicted residual stream derived from the geometric space $f(r,g,p)=(r\cos\theta,r\sin\theta,r)$. In the early layers, errors gradually decrease because the model has not yet captured semantic information, and the geometric space is still being constructed. The continuous decline in error reflects the model’s growing ability to capture semantic information and progressively build a coherent geometric representation. By layer 20, the error stabilizes, indicating that these layers effectively encode the periodic and geometric relationships between atomic properties such as atomic number, group, and period.
Figure B.1: Variation of Absolute Prediction Errors Across Layers with Intervention. The plot shows the mean absolute error (MAE) for each layer, along with the minimum and maximum error range represented by the shaded region. Missing data points were replaced with a value of 50 before computing the absolute errors.
However, beyond layer 30, the error increases sharply as the model begins outputting non-numeric tokens (replaced with an error value of 50 in the plot, corresponding to the maximum possible error given the 50 elements and atomic numbers). This can be attributed to two factors. First, if the numeric token is not the first output, generating the correct answer requires residual streams across all token positions. But only the last residual stream was replaced. Intervening too late disrupts the established flow of residual streams at other positions, which have already determined the output content. Second, it is also likely due to the model shifting its focus from geometric relationships to higher-level abstractions or context-dependent reasoning in the later layers. Therefore, for the intervention experiments on geometric relationships, we selected layer 20 as it balances effective encoding of geometric relationships and minimizes disruption to the model’s output process.
@@ B.2 Detailed evaluation of geometric spaces
The primary evaluation criterion used in the main text is the absolute error threshold ($\leq 2$), as discussed in detail in Sec.3. This metric was chosen because it better captures the accuracy of residual stream interventions. However, other metrics, such as $R^{2}$, Pearson correlation, and qualitative mapping fidelity, also provide valuable insights. These complementary results are summarized in Table B.1.
@@
Table B.1: Performance of different low-dimensional spaces for residual stream intervention. Each space represents a unique pattern, with results assessed using $R^{2}$, Pearson correlation, and percentage of predictions within absolute error $\leq 2$.
@@
In the main paper, we demonstrate two geometric space intervention results; however, other shapes can also be extracted. Fig. B.2 shows the extracted linear structure from interventions. While the alignment of points along a straight path indicates the presence of a linear structure, the overlapping points suggest its limitations in distinguishing atomic number. Compared to more expressive shapes like spirals, linear structures may struggle to effectively capture periodic or distinct features.
@@
Figure B.2: Helix and Linear Structure in the Geometric Space from Intervention Experiments. The figure shows predictions (colored points) and their alignment with the ground truth (gray line).
Appendix C Attention map detailed results
To investigate how the model prioritizes different parts of the input text, we conducted a preliminary analysis using the 32-layer Meta-Llama-3-8B model. We adopted the attribute $A_{j}$, Period and Group, and iterated over $X_{i}$, consisting of 50 elements, using the prompt template: ‘In the periodic table of elements, the $A_{j}$ of $X_{i}$ is.’ These prompts were input into the language model, and we analyzed the average attention across all attention heads in each transformer layer from the token ‘is’ to all other tokens. The averaged results across different prompts are presented in Fig.C.1.
Figure C.1: Average attention distribution analysis of the 32-layer Meta-Llama-3-8B model across transformer layers, based on prompts, ‘In the periodic table of elements, the $A_{j}$ of $X_{i}$ is,’ where $A_{j}$ is an attribute (period or group) and $X_{i}$ is an element. The heatmap (left) shows average attention from ‘is’ to all tokens, while line plots (right) depict attention to target tokens (e.g., element and attribute), average attention to other tokens, and attention entropy. Intermediate layers focus on significant token relationships with higher entropy, while later layers (excluding the final layer) show evenly distributed attention and lower entropy, reflecting a shift to broader context integration.
Appendix D Logit len and tuned len
We input the prompt ‘The atomic number of Mg is ’ and analyzed the token probabilities at each layer using logit-len. By normalizing the final token’s hidden state with LayerNorm and applying the vocabulary head followed by softmax, we obtained the top-ranked tokens directly output by each layer. In each layer, we extracted the probability of the target token, $t_{\text{target}}$ — the output token from the last layer, and checked if it ranked within the top 50 most probable tokens for that layer. The results are shown in Fig. D.1.
Figure D.1: Probability of the target token $t_{\text{target}}$ across layers in Meta-Llama-3-8B for the prompt “The atomic number of Mg is”. Following the logit lens framework, each line shows $t_{\text{target}}$’s probability derived from intermediate layer logits. Probabilities are computed by iteratively re-running the model with the next token added. Markers indicate layers where $t_{\text{target}}$ ranks in the top 50 most probable tokens.
We also performed additional experiments using tuned lens in Fig D.2. The results consistently show a similar trend: tokens such as “of” and “is” achieve high prediction accuracy at intermediate or earlier layers, while more factual or critical token (e.g., blue line “5” only becomes the top prediction at layer 29)—require deeper processing.
Figure D.2: Probability of the target token $t_{\text{target}}$ across layers in Meta-Llama-3-8B for the prompt “The atomic number of Mg is”. Each line shows $t_{\text{target}}$’s probability computed using a tuned lens—a learned linear probe trained to decode intermediate hidden states. Probabilities are obtained by iteratively re-running the model with the next token appended. Markers indicate layers where $t_{\text{target}}$ ranks among the top prediction.
Appendix E Attribute representations overlap in intermediate layers
E.1 Probing weights analysis
As outlined in Sec. 4.1, we trained a linear model for each attribute $A_{j}$ at each layer $l$, yielding a weight vector $\mathbf{w}_{j}^{(l)}$ that represents how attribute $A_{j}$ is stored in the residual stream space of layer $l$. To analyze attribute relationships across layers, we computed the cosine similarity between weight vectors of different attributes using continuation-style residual stream sets to minimize language pattern influence.
Figure E.1: Cosine similarity between weight vectors of linear probes for attribute pairs across layers in Meta-Llama-3.1-70B. The shaded area (99.9% CI) represents unrelated boundaries.
Fig. E.1 illustrates the cosine similarity across 80 layers of Meta-Llama-3.1-70B. Notably, in high-dimensional spaces, random vector pairs typically approach orthogonality due to the ‘blessing of dimensionality’. To illustrate this, we randomly sampled vector pairs in an 8129-dimensional space (the residual stream vector size of Meta-Llama-3.1-70B) and calculated their cosine similarity, with the 99.9% confidence interval (CI) shown in gray. Cosine similarity outside this interval indicates meaningful relationships between attributes. See Appendix E.2 for more details.
E.2 Blessing of dimensionality